. Calculate the number of moles of each compound
a. 5.26 ng NH4C:l (53.5)
b. 12.78µg strontium nitrate: Sr(ClO3)2 :(255)
c. 6.89 mg Li2CrO4 :(130)
1
Expert's answer
2020-12-15T07:52:06-0500

a. Mole = mass/molar mass = (5.26 × 10-9 )/ 53.5 = 98.32 × 10-12 mole

b. Mole = mass / molar mass = (12.78 × 10-6) / 255 = 5.01 × 10-8 mole

c. Mole = mass/ molar mass = (6.89 × 10-3 ) / 130 = 0.000053 mole